INFRARED SPECTRA OF HYDROGEN CLUSTERS SEEDED WITH CARBON DIOXIDE

\underline{A.R.W. McKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

It is now possible to probe cold (<0.5 K) helium clusters in the size range N ∼ 2 to 70 by means of the vibration-rotation spectrum of an embedded infrared chromophore molecule such as CO_{2},^{a} often with atom-by-atom resolution. To some extent, hydrogen clusters can also be studied in this way, as shown by our previous work in which CO,^{b} OCS,^{c} and N_{2}O^{d} were the chromophores. Here we extend the study of hydrogen clusters to the case of CO_{2} as the probe. The symmetry of CO_{2} provides an important difference compared to the other probe molecules. This has the effect of eliminating half of the rotational levels (for the normal C^{16}O_{2} or C^{18}O_{2} isotopomers) and of accentuating the differences between paraH_{2} and orthoH_{2} clusters. As in the case of (H_{2})_{N}-OCS and (H_{2})_{N}-N_{2}O, we find that (H_{2})_{N}-CO_{2} cluster transitions are relatively easy to identify up to about N = 7, but difficult to follow above this point. However, in contrast to the previous work there is intriguing evidence for a series of weak but regularly-spaced transitions which may extend to N ∼ 15 or beyond.

aA.R.W. McKellar, J. Chem. Phys. 128, 044308 (2008).

bS. Moroni, M. Botti, S. De Palo, and A.R.W. McKellar, J. Chem. Phys. 122, 094314 (2005).

cJ. Tang and A.R.W. McKellar, J. Chem. Phys. 121, 3087 (2004).

dJ. Tang and A.R.W. McKellar, J. Chem. Phys. 123, 114314 (2005).